Section 8

Analysis

In case of several authors, Invited Speakers are marked with a ^{∗}.

Kari Astala: Analytic Aspects of Quasiconformality . . . . . . . . . . . . . . . II 617
Michael Christ: Singularity and Regularity — Local and Global … II 627
Nigel Higson: The Baum-Connes Conjecture . . . . . . . . . . . . . . . . . . . . . . . II 637
Michael T. Lacey: On the Bilinear Hilbert Transform . . . . . . . . . . . . . II 647

Pertti Mattila: Rectifiability, Analytic Capacity, and Singular Integrals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . II 657

Vitali Milman: Randomness and Pattern in Convex Geometric Analysis . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . II 665

Detlef Müller: Functional Calculus on Lie Groups and Wave Propagation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . II 679

Stefan Müller^{∗} and Vladimir Šverák: Unexpected Solutions of First and Second Order Partial Differential Equations . . . . . . . . . . . . . . II 691

Klas Diederich and Sergey Pinchuk^{∗}: Reflection Principle in Higher Dimensions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . II 703

Kristian Seip: Developments from Nonharmonic Fourier Series . . . . . II 713
Hart F. Smith: Wave Equations with Low Regularity Coefficients . . II 723

Nicole Tomczak-Jaegermann: From Finite- to Infinite-Dimensional Phenomena in Geometric Functional Analysis on Local and Asymptotic Levels . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . II 731

Stephen Wainger: Discrete Analogues of Singular and Maximal Radon Transforms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . II 743

Thomas Wolff: Maximal Averages and Packing of One Dimensional Sets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . II 755

Documenta Mathematica · Extra Volume ICM 1998 · II ·

616

Documenta Mathematica · Extra Volume ICM 1998 · II ·